The concentration of fructose 6-phosphate (F6P) an
Answer #1

\Delta G = \Delta G^{0'} + RTln \frac{(F-6-P) (ATP)}{(F-1,6-BiP)(ADP)}

\Delta G = -18.3 + (8.314)(310)ln \frac{(0.089)(12.0)}{(0.012)(1.2)}

\Delta G = -18.3 + 2577.34 ln 1.068 / 0.0144

\Delta G = -18.3 + 2577 .34 ln 74.16

\Delta G = -18.3 + 2577 .34 (4.306) = 112079 / 1000 = 11.2079

\Delta G = + 11.2

Option e is correct.